OpenStudy (anonymous):

1. Rationalize the denominator of 5/ square root 3 2. Square root 10 X square root 8 3. Square root 5 ( 8+3 square root 6)

OpenStudy (anonymous):

@Shane_B make sure you read it right haha :)

OpenStudy (shane_b):

Well, the first one is just like the one I posted before...same procedure :P

OpenStudy (shane_b):

Here's the first one:\[(\frac{5}{\sqrt{3}})(\frac{\sqrt{3}}{\sqrt{3}})=\frac{5 \sqrt {3}}{3}\]

OpenStudy (shane_b):

@Aditi_Singh is working so I guess he will post the others

OpenStudy (anonymous):

1. Solve it the way Shane did it in your previous question. 2. use the fact that - \[\sqrt a \times \sqrt b = \sqrt{ab}\]

OpenStudy (anonymous):

haha thanks :>

OpenStudy (anonymous):

@Shane_B : " She" :/

OpenStudy (anonymous):

^ lol true

OpenStudy (shane_b):

I didn't even look at the picture...my fault :)

OpenStudy (anonymous):

wiat im a little confused still on the 2nd and 3rd question guys o.O

OpenStudy (anonymous):

3.\[\sqrt5 ( 8+3 \sqrt 6)\] \[= 8\sqrt5 + 3\sqrt30\]

OpenStudy (shane_b):

What SHE means for #2 is: \[(\sqrt{10})(\sqrt{8})=\sqrt{8*10}=\sqrt{80}\]

OpenStudy (anonymous):

@Shane_B : Lol :P :D

OpenStudy (anonymous):

thanks for the help on #2! :) but im still lost on #3 o.o

OpenStudy (shane_b):

I'm serious think I'm having a bout of attention deficit disorder at the moment...need coffee.

OpenStudy (anonymous):

^lol first the reading confusion, then gender confusion o.O

OpenStudy (anonymous):

guys I need help on #3 pleasee

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I solved it for you above .. is it wrong ?? :P :/

OpenStudy (anonymous):

OH nvm just saw that lol thankss (now im reading things wrong o.o)

OpenStudy (shane_b):

HAH!

OpenStudy (anonymous):

@Shane_B you're giving me your reading disability o.o

OpenStudy (anonymous):

Hahahaha :P .. literally, ROFL! :P